OpenStudy (anonymous):

My teacher says I still have the wrong answer! I need your help. It costs XYZ Corporation $7500 to produce 1400 units and $4440 to produce 500 units. Let C represent the cost in dollars and x represent the number of units produced. Find a linear equation for C in terms of x.

OpenStudy (mertsj):

Your ordered pairs are (1400,7500) and (500,4440) Find the slope and y intercept using those points.

OpenStudy (mertsj):

m=3060/900=17/5

OpenStudy (anonymous):

C = y , Number of units produced = x

OpenStudy (mertsj):